# Scale-Equivariant Imaging: Self-Supervised Learning for Image Super-Resolution and Deblurring

### v3

@@ I Introduction
where $y$ represents the observed measurements, $z$ represents the latent image, $A$ is the imaging operator, and $\varepsilon$ depicts the noise corrupting the measurements. In this paper, we focus on additive white Gaussian noise but our results can easily be generalized for other noise distributions, e.g., Poisson, or mixed Gaussian-Poisson. Recovering $z$ from the noisy measurements is generally ill-posed, as the forward operator is often rank-deficient, e.g., when the number of measurements is smaller than the number of image pixels. Moreover, even when the forward operator is not rank-deficient, it might still be ill-conditioned, and reconstructions might be very sensible to noise (for example, in image deblurring problems).
@@
In recent years, self-supervised methods have highlighted the possibility of learning from measurement data alone, thus bypassing the need for ground truth references. These methods focus either on taking care of noisy data and/or a rank-deficient operator. Noise can be handled through the use of Stein’s unbiased risk estimator if the noise distribution is known [4], or via Noise2x techniques [5, 6, 7] which require milder assumptions on the noise distribution such as independence across pixels [7]. Rank-deficient operators can be handled using measurement data associated with multiple measurement operators $A_{1},\dots,A_{G}$ (each with a different nullspace) [8][9], or by assuming that the image distribution is invariant to certain groups of transformations such as translations or rotations [10, 11].
Self-supervised image deblurring methods have been proposed in settings where the images are blurred with multiple kernels [12], however this assumption is not verified in many applications. Moreover, self-supervised methods that rely on a single operator (e.g., a single blur kernel) [10, 11], have not been applied to inverse problems with a single forward operator that removes high-frequency information such as image super-resolution and image deblurring, except for methods exploiting internal/external patch recurrence [13] which assume noiseless measurements, thus not being easily generalizable to other inverse problems.
Image deblurring and super-resolution are two inverse problems particularly difficult to solve using self-supervised methods. The former is severely ill-conditioned, and the latter is rank-deficient. This makes methods that minimize the measurement consistency error inapplicable as they cannot retrieve any information that lies within the nullspace [10] or where the forward operator is ill-conditioned, whenever there is even the slightest amount of noise present.
@@
Figure 1: Scale-equivariant imaging. The proposed loss is the sum of the SURE loss $\mathcal{L}_{\text{SURE}}(\theta)$ which penalizes reconstruction error in the measurement domain, and of the scale-equivariant loss $\mathcal{L}_{\text{SEQ}}(\theta)$ by viewing $x^{(1)}=f_{\theta}(y)$ as a ground truth which might lack high-frequency content, and the downscaled image $x^{(2)}$ as another one which has high-frequency content, and finally by penalizing the mean squared error of the reconstruction $x^{(3)}$ obtained from a noisy measurement of $x^{(2)}$. The difference in high-frequency content is shown by $\hat{x}^{(1)}$ and $\hat{x}^{(2)}$, the discrete Fourier transforms of $x^{(1)}$ and $x^{(2)}$. One of our key contributions is stopping the gradient of $x^{(2)}$ during the stochastic gradient descent.
@@ II Related work
A family of super-resolution and image deblurring methods rely on plug-and-play priors [19] and optimization methods. Other approaches rely on the inductive bias of certain convolutional architectures, i.e., the deep image prior [20, 21]. Some methods combine both techniques [22]. However, these approaches can provide sub-optimal results if the prior is insufficiently adapted to the problem.
@@
The equivariant imaging (EI) method leverages the invariance of typical image distributions to rotations and/or translations to learn from measurement data alone, in problems where the forward operator is incomplete such as sparse-angle tomography, image inpainting and magnetic resonance imaging [10, 11, 26]. However, this approach requires the forward operator to not be equivariant to the rotations or translations. The super-resolution problems and image deblurring problems with an isotropic blur kernel do not verify this condition, and existing EI methods cannot be used.
@@
Internal patch recurrence of natural images can be exploited to learn to super-resolve images without high-resolution references [13]. However, while this self-supervised learning method can be extended to collections of low-resolution images via external patch recurrence, it cannot be easily generalized to more general inverse problems (e.g., image deblurring) and noisy measurement data. The method proposed in this work generalizes the idea of patch recurrence to scale-invariance and can be applied to any imaging inverse problem where high-frequency information is missing.
@@ III Learning from bandlimited images
We consider images as functions $z:\mathbb{R}^{2}\to\mathbb{C}^{d}$ where $d\geq 1$ denotes the number of channels. Formally, we model the set of images as a subset $\mathcal{Z}$ of the Schwartz space [^1] The Schwartz space $\mathcal{S}$ is dense in $L^{2}$, and unlike $L^{2}$ it is closed under convolution, making it a very general model for studying bandlimited inverse problems. $\mathcal{S}$ of functions mapping $\mathbb{R}^{2}$ to $\mathbb{C}^{d}$. The problems of super-resolution and deblurring fall under the umbrella of image recovery from bandlimited measurements. Thus in this section, to simplify notation we assume that the forward operator is modelled as $y={h}*z$, where ${h}\in\mathcal{S}$ is a filter acting on images $z$ by convolution. Moreover, we assume that ${h}$ has a finite bandwidth $0\leq\xi_{h}<\infty$, defined as:
@@
While physical images $z$ are typically not bandlimited, the measurements $y={h}*z$ have bandwidth at most $\xi_{{h}}$, and due the Shannon-Nyquist’s sampling theorem, they can be represented by discrete samples on a regular grid of sampling rate $2\xi_{{h}}$ [29].
@@
Figure 2: Spectral effect of rescaling. In bandlimited problems such as image super-resolution and deblurring, the image frequencies are only observed on a low-frequency band $[0,\xi_{h})$ determined by the bandlimiting operator (e.g., blur and anti-aliasing filters), and the goal is to recover texture information in a higher frequency band $(\xi_{h},\xi_{\varphi})$ where $\xi_{\varphi}/\xi_{h}$ can be interpreted as the increase in effective resolution. Scale-invariance makes this possible as the spectral information is available at all scales: an image $z({u})$ and the rescaled image $\Sigma_{s}z({u})$ with scaling factor $s$ have the same spectral content except in different frequency bands.
@@
Let ${h},\varphi\in\mathcal{S}$. If ${h}$ and $\varphi$ are bandlimited with bandwidths $\xi_{h}<\xi_{\varphi}$, then there are sets $\mathcal{Z}_{1},\mathcal{Z}_{2}\subseteq\mathcal{S}$ such that
@@
$\varphi$ is bandlimited, and $\hat{h}(0)\neq 0$, where $\hat{h}$ is the Fourier transform of ${h}$, then the set $\mathcal{X}$ defined in Eq. 4 is uniquely determined by the set of observed images $\mathcal{Y}$ defined in Eq. 3. More precisely, there is a $s\geq\frac{\xi_{\varphi}}{\xi_{h}}$ which only depends on $\varphi$ and ${h}$, for which $\hat{h}(s^{-1}\xi)\neq 0$ as long as $\lVert\xi\rVert_{2}<\xi_{\varphi}$, and
@@ IV Proposed method
In this section, we assume that there is an approximately one-to-one correspondence between the latent images and the measured images which guarantees that image recovery is possible, and we expose our proposed method for deblurring and super-resolution that exploits the scale-invariance of typical image distributions.
@@
where $\downarrow_{r}$ is the subsampling operator by factor $r$, and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is additive Gaussian noise [^2] We focus on Gaussian noise in this work and use the corresponding risk estimator – it is meant to be substituted for other noise distributions.. In super-resolution problems ${h}$ is the anti-aliasing filter and $r>1$ whereas in deblurring problems ${h}$ is the blur kernel and $r=1$. Self-supervised learning-based methods attempt to solve this problem by learning a reconstruction function [^3] For the problem to be well-defined, it is necessary that there is one-to-one-ness between the large-bandwidth images and the small-bandwidth images. $f_{\theta}:y\mapsto x$ parametrized by a neural network with learnable weights $\theta\in\mathbb{R}^{p}$ using a dataset of blurry/low-resolution images only $\{y_{i}\}_{i=1}^{N}$.
@@
In the forward pass, we first compute $x_{i,\theta}^{(1)}=f_{\theta}(y_{i})$, which we view as an estimation of the $i$th ground truth image $x_{i}$. Next, we compute $x_{i}^{(2)}=\mathop{\Sigma_{s}}x_{i,\theta}^{(1)}$, where $\Sigma_{s}$ is the downscaling transform by factor $s\in(0,1)$, without propagating the gradient, which effectively leaves $\nabla_{\theta}x_{i}^{(2)}=0$. Finally, we compute $x_{i,\theta}^{(3)}=f_{\theta}(Ax_{i}^{(2)}+\tilde{\varepsilon})$ where $\tilde{\varepsilon}$ is sampled from $\mathcal{N}(0,\sigma^{2}I)$. The three terms are used to compute $\mathcal{L}_{\text{SURE}}$ and $\mathcal{L}_{\text{SEQ}}$, where $\mathcal{L}_{\text{SEQ}}$ is defined as
@@
[…]
@@
Downsampling (×$\times 2$) Downsampling (×$\times 3$) Downsampling (×$\times 4$)
@@
The CSS method requires that the forward operator is a downsampling transformation and that the noise $\varepsilon$ is sufficiently small, such that $\left({h}*y\right){\downarrow_{r}}\approx x^{\prime}$ for some $x^{\prime}$ which also belongs to the dataset. On the contrary, as the proposed method exploits scale-invariance on the reconstructed images, it can be used with any operator $A$. Thus, for the super-resolution problems with a small amount of noise, we expect to obtain a similar performance with both methods, whereas in the more challening image deblurring problem we expect to observe a different performance across methods.
@@
Gaussian filter (medium) Box filter (medium) Downsampling (×$\times 2$)
@@ V Experiments
Figure 5: Point-spread functions. Our method makes no assumption about the shape or strength of the blur kernel making it broadly applicable and in particular we verify that it performs well on Gaussian filters with standard deviation $1$, $2$ or $3$ pixels, and for box filters with radius $2$, $3$ or $4$ pixels.
For every degradation model, we constitute pairs of clean and corrupted images that we use for training and evaluation. We evaluate all methods, using two dataset of natural images DIV2K [37] and Urban100 [38], and a dataset of computed tomography (CT) scans [39]. Figure 5 shows the 6 different blur kernels used for synthesizing some of the corrupted images, and the remaining ones are synthesized using bicubic downsampling with multiple sample rates (2 and 4). The resulting images are further corrupted by additive white Gaussian noise making the degradation model closer to real-world image degradations happening in most imaging systems. The standard deviation for the Gaussian noise is set to $\sigma=5/255$. The 900 image pairs from DIV2K are split into 100 testing pairs and 800 training pairs, the 100 image pairs from Urban100 are split into 10 testing pairs and 90 testing pairs, and the 5092 pairs for CT scans are split into 100 testing pairs and 4992 training pairs. Tables I and IV show that scale-invariant imaging tends to perform as well as the supervised baseline and significantly better than the other state-of-the-art self-supervised methods, both across different imaging modalities and different image distributions. Examples of deblurred and upsampled images are shown in Figure 4.
@@
Because it is fully self-supervised and unlike supervised methods, it is possible to use scale-equivariant imaging directly on test data. This enables fine-tuning models pre-trained on image distributions that might differ from the test data, generally resulting in better performant models. We evaluate this approach by fine-tuning 3 models on CT scans. The 3 models are pre-trained on pairs high and low-resolution images from DIV2K (natural images), each for a different sampling rate (2, 3, and 4), all with the same supervised training algorithm. The low-resolution images are synthesized using bicubic downsampling similarly to the low-resolution CT scans that we are using except for the absence of additional noise. Table III shows that the performance of the fine-tuned model is significantly greater than the performance of the pre-trained model, ranging from a .5.5.5.5 dB to a $2$ dB difference in PSNR. Example of upsampled images are shown in Figure 6.
@@
In addition to the experiments comparing our method with various other methods, we also conduct an ablation study to evaluate the importance of gradient stopping, one of the key ingredients in scale-equivariant imaging. For each of the three degradation models, we train two models using our proposed method: one with and one without gradient stopping. Table II shows that gradient stopping improves the reconstructions by more than $2$dB (PSNR) in all three cases. This is consistent with our broader empirical observation that gradient stopping consistently improves the performance of the proposed method.
@@
Downsampling (×$\times 2$) Downsampling (×$\times 3$) Downsampling (×$\times 4$)

### v4

@@ I Introduction
where $y$ represents the observed measurements, $z$ represents the latent image, $A$ is the imaging operator, and $\varepsilon$ depicts the noise corrupting the measurements. In this paper, we focus on additive white Gaussian noise but our results can easily be generalized for other noise distributions, e.g., Poisson, or mixed Gaussian-Poisson. Recovering $z$ from the noisy measurements is generally ill-posed, as the forward operator is often rank-deficient, e.g., when the number of measurements is smaller than the number of image pixels. Moreover, even when the forward operator is not rank-deficient, it might still be ill-conditioned, and reconstructions might be very sensitive to noise (for example, in image deblurring problems).
@@
In recent years, self-supervised methods have highlighted the possibility of learning from measurement data alone, thus bypassing the need for ground truth references. These methods focus either on taking care of noisy data and/or a rank-deficient operator. Noise can be handled through the use of Stein’s unbiased risk estimator if the noise distribution is known [4], or via Noise2x techniques [5, 6, 7] which require milder assumptions on the noise distribution such as independence across pixels [7]. Rank-deficient operators can be handled using measurement data associated with multiple measurement operators $A_{1},\dots,A_{G}$ (each with a different nullspace) [8][9], or by assuming that the image distribution is invariant to certain groups of transformations such as translations or rotations [10, 11]. This assumption is generally valid as a shifted or rotated image is generally still a valid image.
Self-supervised image deblurring methods have been proposed in settings where the images are blurred with multiple kernels [12], however this assumption is not verified in many applications. Moreover, self-supervised methods that rely on a single operator (e.g., a single blur kernel) [10, 11], have not been applied to inverse problems with a single forward operator that removes high-frequency information such as image super-resolution and image deblurring, except for methods exploiting internal/external patch recurrence [13] which assume noiseless measurements, and thus do not easily generalize to other inverse problems.
Image deblurring and super-resolution are two inverse problems that are particularly difficult to solve using self-supervised methods. The former is severely ill-conditioned, and the latter is rank-deficient. This makes methods that minimize the measurement consistency error inapplicable as they cannot retrieve any information that lies within the nullspace [10] or where the forward operator is ill-conditioned, whenever there is even the slightest amount of noise present.
@@
Figure 1: Overview of the method. Our scale-equivariant imaging loss $\mathcal{L}_{\mathrm{SEI}}$ combines a measurement consistency loss that is robust to noise (SURE) with a new self-supervised loss that substitutes downscaled reconstructions for ground truth images. The use of downscaling widens the frequency spectrum past the cut-off imposed by the blurring and super-resolution operators, thereby enabling the recovery of high-frequency information lost in the measurement process.
@@ II Related work
A family of super-resolution and image deblurring methods rely on plug-and-play or diffusion priors [19, 20, 21] and optimization methods. Other approaches rely on the inductive bias of certain convolutional architectures, i.e., the deep image prior [22, 23]. Some methods combine both techniques [24]. However, these approaches can provide sub-optimal results if the prior is insufficiently adapted to the problem.
@@
Typical image distributions $\mathcal{Z}$ are invariant to rotations and/or translations, i.e., $\mathop{T_{g}}z\in\mathcal{Z},\ \forall z\in\mathcal{Z}$, for every roto-translation $\{T_{g}\}_{g=1}^{|G|}$. Equivariant imaging (EI) leverages this property to learn reconstruction functions from measurement data alone, in problems where the forward operator is incomplete such as sparse-angle tomography, image inpainting, image fusion and magnetic resonance imaging [10, 11, 28, 29]. However, this approach requires the forward operator not to be equivariant to rotations or translations, i.e., $A(\mathop{T_{g}}z)\neq\mathop{T_{g}}A(z)$ for at least some $z$ and $T_{g}$. The super-resolution and deblurring problems generally do not verify this condition, and existing EI methods cannot be used. The identification of invariant image distributions has been studied in a subsequent work [9], but only for discrete images under the action of a compact group, with guarantees for almost every generic operator. We introduce a different analysis imposed by the non-compact scaling group which operates on continuous images. This new framework leverages Fourier analysis instead of representation theory and yields guarantees of a stronger kind, covering every bandlimiting operator as opposed to almost every.
@@
Internal patch recurrence of natural images can be exploited to learn to super-resolve images without high-resolution references [13]. However, while this self-supervised learning method can be extended to collections of low-resolution images via external patch recurrence, it cannot be easily generalized to other inverse problems where the degraded image has different statistics from the ground truth image (e.g., image deblurring and noisy measurement data). The method proposed in this work generalizes the idea of patch recurrence to scale-invariance and can be applied to any imaging inverse problem where high-frequency information is missing.
@@ III Learning from bandlimited images
We consider images as functions $z:\mathbb{R}^{2}\to\mathbb{C}^{d}$ where $d\geq 1$ denotes the number of channels. Formally, we model the set of images as a subset $\mathcal{Z}$ of the Schwartz space [^1] The Schwartz space $\mathcal{S}$ is dense in $L^{2}$, and unlike $L^{2}$ it is closed under convolution, making it a very general model for studying bandlimited inverse problems. $\mathcal{S}$ of functions mapping $\mathbb{R}^{2}$ to $\mathbb{C}^{d}$. The problems of super-resolution and deblurring fall under the umbrella of image recovery from bandlimited measurements. Thus in this section, to simplify notation we assume that the forward operator is modelled as y=hzy={h}*z, where ${h}\in\mathcal{S}$ is a filter acting on images $z$ by convolution. Moreover, we assume that h{h} has a finite bandwidth $0\leq\xi_{h}<\infty$, defined as:
@@
While physical images $z$ are typically not bandlimited, the measurements y=hzy={h}*z have bandwidth at most $\xi_{{h}}$, and due the Shannon-Nyquist’s sampling theorem, they can be represented by discrete samples on a regular grid of sampling rate $2\xi_{{h}}$ [31].
@@
Figure 2: Spectral effect of rescaling. In bandlimited problems such as image super-resolution and deblurring, the image frequencies are only observed on a low-frequency band $0,\xi_{h})$ determined by the bandlimiting operator (e.g., blur and anti-aliasing filters), and the goal is to recover texture information in a higher frequency band $(\xi_{h},\xi_{\varphi})$ where $\xi_{\varphi}/\xi_{h}$ can be interpreted as the increase in effective resolution. Scale-invariance makes this possible as the spectral information is available at all scales: an image z(u)z({u}) and the rescaled image $\Sigma_{s}z({u})$ with scaling factor $s$ have the same spectral content except in different frequency bands.
@@
Let ${h},\varphi\in\mathcal{S}$. If h{h} and $\varphi$ are bandlimited with bandwidths $\xi_{h}<\xi_{\varphi}$, then there are sets $\mathcal{Z}_{1},\mathcal{Z}_{2}\subseteq\mathcal{S}$ such that
@@
$\varphi$ is bandlimited, and $\hat{h}(0)\neq 0$, where $\hat{h}$ is the Fourier transform of h{h}, then the set $\mathcal{X}$ defined in Eq. 4 is uniquely determined by the set of observed images $\mathcal{Y}$ defined in Eq. 3. More precisely, there is a $s\geq\frac{\xi_{\varphi}}{\xi_{h}}$ which only depends on $\varphi$ and h{h}, for which $\hat{h}(s^{-1}\xi)\neq 0$ as long as $\lVert\xi\rVert_{2}<\xi_{\varphi}$, and
@@ IV Proposed method
In this section, we assume that there is an approximately one-to-one correspondence between the latent images and the measured images which guarantees that image recovery is possible, and we introduce our proposed method for deblurring and super-resolution that exploits the scale-invariance of typical image distributions.
@@
where $\downarrow_{r}$ is the subsampling operator by factor $r$, and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is additive Gaussian noise [^2] We focus on Gaussian noise in this work and use the corresponding risk estimator – it can be substituted for other noise distributions.. In super-resolution problems h{h} is the anti-aliasing filter and $r>1$ whereas in deblurring problems h{h} is the blur kernel and $r=1$. Self-supervised learning-based methods attempt to solve this problem by learning a reconstruction function [^3] For the problem to be well-defined, it is necessary that there is one-to-one-ness between the large-bandwidth images and the small-bandwidth images. $f_{\theta}:y\mapsto x$ parametrized by a neural network with learnable weights $\theta\in\mathbb{R}^{p}$ using a dataset of blurry/low-resolution images only $\{y_{i}\}_{i=1}^{N}$.
@@
[…]
@@
In all experiments, $\Sigma_{s}$ consists in resampling the input image viewed as periodic on a (coarse) grid with a spacing of $s^{-1}$ pixels, using bicubic interpolation. The grid center is chosen randomly at each call and the scale factor is chosen uniformly from $\{.5,.75\}$. Figure 3 shows two possible downscalings of an image patch.
@@
Downsampling ($\times 2$) Downsampling ($\times 3$) Downsampling ($\times 4$)
@@
The CSS method requires that the forward operator is a downsampling transformation and that the noise $\varepsilon$ is sufficiently small, such that $\left({h}*y\right){\downarrow_{r}}\approx x^{\prime}$ for some $x^{\prime}$ which also belongs to the dataset. On the contrary, as the proposed method exploits scale-invariance on the reconstructed images, it can be used with any operator $A$. Thus, for the super-resolution problems with a small amount of noise, we expect to obtain a similar performance with both methods, whereas in the more challenging image deblurring problem we expect to observe a different performance across methods.
@@
Gaussian filter (medium) Box filter (medium) Downsampling ($\times 2$)
@@ V Experiments
Figure 5: Point-spread functions. Our method makes no assumption about the shape or strength of the blur kernel making it broadly applicable and in particular we verify that it performs well on Gaussian filters with standard deviation 11, 22 or 33 pixels, and for box filters with radius 22, 33 or 44 pixels.
For every degradation model, we constitute pairs of clean and corrupted images that we use for training and evaluation. We evaluate all methods, using two datasets of natural images DIV2K [38] and Urban100 [39], and a dataset of computed tomography (CT) scans [40]. Figure 5 shows the 6 different blur kernels used for synthesizing some of the corrupted images, and the remaining ones are synthesized using bicubic downsampling with multiple sampling rates (2 and 4). The resulting images are further corrupted by additive white Gaussian noise making the degradation model closer to real-world image degradations happening in most imaging systems. The standard deviation for the Gaussian noise is set to $\sigma=5/255$. The 900 image pairs from DIV2K are split into 100 testing pairs and 800 training pairs, the 100 image pairs from Urban100 are split into 10 testing pairs and 90 training pairs, and the 5092 pairs for CT scans are split into 100 testing pairs and 4992 training pairs. Tables I and IV show that scale-invariant imaging tends to perform as well as the supervised baseline and significantly better than the other state-of-the-art self-supervised methods, both across different imaging modalities and different image distributions. Examples of deblurred and upsampled images are shown in Figure 4.
@@
Because it is fully self-supervised and unlike supervised methods, it is possible to use scale-equivariant imaging directly on test data. This enables fine-tuning models pre-trained on image distributions that might differ from the test data, generally resulting in better performing models. We evaluate this approach by fine-tuning 3 models on CT scans. The 3 models are pre-trained on pairs high and low-resolution images from DIV2K (natural images), each for a different sampling rate (2, 3, and 4), all with the same supervised training algorithm. The low-resolution images are synthesized using bicubic downsampling similarly to the low-resolution CT scans that we are using except for the absence of additional noise. Table III shows that the performance of the fine-tuned model is significantly greater than the performance of the pre-trained model, ranging from a .5.5 dB to a $2$ dB difference in PSNR. Example of upsampled images are shown in Figure 6.
@@
In addition to the experiments comparing our method with other methods, we also conduct a range of ablation studies to explore the role of the gradient stopping, the sensitivity of key parameters, and the influence of the chosen network architecture and interpolation strategy. Here, we will focus on the results for the gradient stopping and relegate the additional studies to Section -A. For each of the three degradation models, we train two models using our proposed method: one with and one without gradient stopping. Table II shows that gradient stopping improves the reconstructions by more than $2$ dB (PSNR) in all three cases. This is consistent with our broader empirical observation that gradient stopping consistently improves the performance of the proposed method.
@@
Downsampling ($\times 2$) Downsampling ($\times 3$) Downsampling ($\times 4$)